OpenStudy (anonymous):

tiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtired

OpenStudy (mayankdevnani):

\[\huge \sqrt{-72}\] First do factorization of 72

OpenStudy (mayankdevnani):

no!! its wrong

OpenStudy (mayankdevnani):

because 72 is negative and i think it is a imaginary number(i)

OpenStudy (mayankdevnani):

welcome:)

OpenStudy (mayankdevnani):

hey!! i don't know much about complex numbers

OpenStudy (kenljw):

sqrt(-|x|) = isqrt(|x| basic sqrt(-1) =i , i^2 = -1, i^3 = -i ect

OpenStudy (mayankdevnani):

I think \(\large i=\color{green}{IOTA}=\color{red}{imaginary \space number}\)